FOURIER TRANSFORM FAR-INFRARED SPECTROSCOPY OF HN_{2}^{+} ON THE AILES BEAMLINE OF SYNCHROTRON SOLEIL

\underline{O. PIRALI}^{a}, S. GRUET^{a}, M. VERVLOET, Ligne AILES, Synchrotron SOLEIL, L’Orme des Merisiers Saint-Aubin, 91192 Gif sur Yvette Cedex - France.

We report the pure rotation spectrum of HN_{2}^{+} measured by Fourier Transform (FT) absorption spectroscopy in the 20-40 cm^{−1} spectral range. The cations are produced in a liquid nitrogen cooled hollow cathode discharge cell developed on the AILES beamline of synchrotron SOLEIL. The set-up was optimized by recording rotation-vibration spectra of H_{3}^{+}(ν_{2} band centered at 2521 cm^{−1}), HN_{2}^{+} (ν_{1} band centered at 3234 cm^{−1}) and HCO^{+}(ν_{1} band centered at 3089 cm^{−1}). Many rotation-vibration lines have been assigned for each ion and 5 pure rotational transitions have been detected for HN_{2}^{+}. These results demonstrate the feasibility to record far-infrared (far-IR) spectra of cationic species using FT broad band spectroscopy associated to the bright synchrotron radiation continuum as an alternative to laser-based frequency tunable techniques. In the presentation, we will describe the experimental set-up, the results obtained and the perspectives of this preliminary work which will rely on the exploitation of intense Coherent Synchrotron Radiation (CSR).

aAlso at: Institut des Sciences Moléculaires d’Orsay, UMR 8214 CNRS-Université Paris-Sud, Bât. 210, 91405 Orsay cedex, France.